Find the required sample size needed to make a 99% confidence interval for the population proportion if the margin of error can be no more than 5%.

  1. Assume you may use

Assume you have no information about .
Answer #1

Here we have: d=1-0.99 = 0.01 (99confidence level) E-0.05 Critical value: Z 4 2 = 2000/2 = 2.576 sample size = E 2012 pipilin